OpenStudy (anonymous):

solve for X: 3/x=4/8

OpenStudy (anonymous):

4/8 is the same as 1/2. Your equation is now 3/x=1/2. Multiply both sides by x to get x into the numerator. You now have 3=x/2. Then multiply both sides of the equation by 2, which gives you 6=x. Problem solved.